OpenStudy (anonymous):

8m^2-15=14m I know I am making this quadratic, but what do I do with the 8??

OpenStudy (gamer56):

Is it ok if i show you all of the steps?

OpenStudy (gamer56):

This is your answer and it will also show you what to do with the 8. Simplifying 8m2 + -15 = 14m Reorder the terms: -15 + 8m2 = 14m Solving -15 + 8m2 = 14m Solving for variable 'm'. Reorder the terms: -15 + -14m + 8m2 = 14m + -14m Combine like terms: 14m + -14m = 0 -15 + -14m + 8m2 = 0 Factor a trinomial. (-3 + -4m)(5 + -2m) = 0

OpenStudy (anonymous):

I don't understand the trinomial part

OpenStudy (triciaal):